OpenStudy (anonymous):

Which equation represents the line passing through the points (1, 4) and (-2, 13)?

OpenStudy (ajprincess):

Equation of the line=> (y-y1)/(x-x1)=(y1-y2)/(x1-x2) (y-4)/(x-1)=(4-13)/(1-(-2)) (y-4)/(x-4)=-9/3 (y-4)/(x-4)=-3 y-4=-3(x-4) See if u cn simplify to get the eqn.